# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ 6.1 Toy Example
and $u(\theta)=[\cos\theta,\sin\theta,0,\dots,0]^{\top}\in\operatorname{\mathbb{R}}^{d}$. The solution set is the $(\texttt{k}-1)$-dimensional sphere of radius $\sqrt{1+\theta^{2}}$ embedded in $\mathbb{R}^{d}$ along the first k coordinates:
@@
    V-PBGD (shen2025penalty): joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\gamma(g(\theta,x)-g^{*}(\theta))$,
@@
    G-PBGD (shen2025penalty): joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\tfrac{\gamma}{2}\|\nabla_{x}g(\theta,x)\|^{2}$,
@@
    RHG (franceschi2017forward): truncated reverse‐mode hypergradient with re‐initialization of $\theta$ every outer step,
@@
    IAPTT-GM (liu2021towards): unroll $T$ inner steps $x_{k+1}=x_{k}-\beta\nabla_{x}g(\theta,x_{k})$ from $x_{0}=z$, pick the pessimistic truncation $k_{\star}=\arg\max_{k}F(\theta,x_{k})$, and update $\theta,z$ by gradient descent on the truncated objective $F(\theta,x_{k_{\star}})$.
All methods are implemented in JAX with identical initialization $\theta_{0}=1.0$, $x_{0}\sim\mathcal{N}(0,I_{d})$, and random seed. We sweep dimension $d\in\{2,5,10,20\}$.
@@
Figure 1 plots $F(\theta_{n})-F^{*}$ over iterations for $d=2$. The zeroth-order method rapidly converges toward zero, while the other methods tend to plateau or oscillate around suboptimal objective values, primarily due to the absence of a minima selection phase.
@@ 6.1.1 Dimension sweep and minima‐selection
5 0.273 0.160 0.302 0.151 $\mathbf{0.015\pm 0.001}$
10 0.169 0.153 0.303 0.152 $\mathbf{0.023\pm 0.001}$
20 0.113 0.174 0.302 0.154 $\mathbf{0.043\pm 0.002}$
Table 1: Final absolute error $|\theta_{N}-\theta^{*}|$ after $N=500$ iterations, evaluated across different dimensions $d$. The manifold dimension $\dim(\mathcal{S}(\theta))$ is d1d-1.
In this experiment, we examine the setting where the intrinsic dimension scales with the ambient dimension, specifically setting $\texttt{k}=d-1$. This simplifies the lower-level objective to $g(\theta,x)=\frac{1}{4}(\|x\|^{2}-\theta^{2})^{2}-\frac{1}{2}\|x\|^{2}$. Table˜1 reports the final absolute error $|\theta_{N}-\theta^{*}|$ after $N=500$ iterations across varying dimensions $d$. The influence of the manifold dimension $\dim(\mathcal{S}(\theta))=d-1$ on the absolute error is also evident. Despite this scaling, our algorithm consistently outperforms the three baseline methods, achieving a significantly closer approximation to the global minimizer $\theta^{\ast}=0$. For robustness, we report the mean and confidence intervals over $10$ independent runs for our randomized method, compared to single runs for the deterministic baselines.
@@ 6.1.2 Constant manifold dimension
We next investigate the distinction between intrinsic and ambient dimensionality by fixing the manifold dimension at $\texttt{k}=2$ while varying the ambient dimension $d$. To align the implementation with our theoretical bounds, we chose the number of Langevin Monte Carlo (LMC) iterations and the batch size $M$ to closely approximate the target Gibbs distribution. As demonstrated in Table˜2, the final error of Algorithm˜2 remains nearly constant as $d$ increases, confirming that the complexity is almost independent of the ambient dimension.
@@
5 0.495 0.313 0.302 0.119 $\mathbf{0.015\pm 0.001}$
10 0.499 0.313 0.316 0.114 $\mathbf{0.018\pm 0.001}$
20 0.511 0.315 0.307 0.121 $\mathbf{0.018\pm 0.002}$
30 0.532 0.313 0.311 0.121 $\mathbf{0.02\pm 0.002}$
Table 2: Final absolute error $|\theta_{N}-\theta^{*}|$ after $N=500$ iterations, evaluated across different dimensions $d$. The manifold dimension $\dim(\mathcal{S}(\theta))=1$ remains constant. Additionally, Algorithm˜2 achieves improved accuracy when using a fixed batch size of $M=400$.
@@
5 $\mathbf{0.011\pm 0.001}$ $\mathbf{0.021\pm 0.001}$
10 $\mathbf{0.025\pm 0.002}$ $\mathbf{0.023\pm 0.002}$
20 $\mathbf{0.030\pm 0.002}$ $\mathbf{0.026\pm 0.002}$
30 $\mathbf{0.035\pm 0.003}$ $\mathbf{0.027\pm 0.002}$
Table 3: Final absolute error $|\theta_{N}-\theta^{*}|$ after $N=500$ iterations, evaluated over lower-level dimension $\delta$ using Algorithm˜2 in the pessimistic setting. The first column corresponds to the pessimistic formulation (33) with lower-level $g_{d-1}$ from (32); the second uses the same problem with lower-level $g_{1}$ from (32).
@@ 6.2 Data hyper-cleaning
In this section, we evaluate our method on the data hyper-cleaning benchmark (franceschi2017forward; shaban2019truncated), a canonical BLO task widely used in the machine learning community. Given a set of corrupted [^19] A sample is clean if its label equals the true class; otherwise it is corrupted. In MNIST, we corrupt by flipping the label of a fraction $p$ of training points uniformly to a different (incorrect) class. $p$ is the pollute-rate. training samples $\mathcal{D}_{\text{tr}}=\{(d_{i}^{\text{tr}},\,\ell_{i}^{\text{tr}})\}_{i=1}^{N}$, a small clean validation set $\mathcal{D}_{\text{val}}=\{(d_{i}^{\text{val}},\,\ell_{i}^{\text{val}})\}_{i=1}^{M}$, and a clean test set, the goal is to learn a data cleaner that down-weights mislabeled training points so as to improve generalization on unseen data. Let $\theta$ denote the data-cleaner parameters (producing per-example weights) and $x$ the model parameters. We use a sigmoid parameterization for the weights, $\omega_{i}(\theta)\;=\;(1+\exp(-x\theta_{i}))^{-1},\ (i=1,\dots,N),$ and constrain $\|\theta\|\leq B$ to avoid exploding weights. The task is cast as the optimistic bilevel problem in the following:
@@
We use MNIST with N=5, [^0] N=5{,}000 training, $M=50$ validation, and 10,00010{,}000 test examples, and uniformly corrupt a fraction $\rho\in\{0.4,0.6,0.8\}$ of the training labels. In practice, the corrupted/imprecise data are widely available, but clean/precise data are scarce. Thus we intentionally adopt a very small validation set. Much of the literature uses roughly equal training/validation splits, in which case the effect of the corruption rate in the uncleaned training data during training is often not apparent. We compare our zeroth-order method (Algorithm˜2) against V-PBGD and G-PBGD (shen2025penalty), RHG (franceschi2017forward), and IAPTT-GM (liu2021towards), using both a linear model and a two-layer MLP which are standard Neural network models in data hyper-cleaning tasks.
@@ Results.
Table˜4 reports solution quality (test accuracy and F1 score [^20] F1-score is defined as follows: We take clean samples as the positive class and corrupted samples as the negative class. Each sample $i$ is assigned a score $s_{i}=\sigma(\theta_{i})$. We predict “clean” for the top (1p)(1-p)-fraction by $s_{i}$. We compute $\text{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\ \text{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$ where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{FN}$ denote true positives, false positives, and false negatives, respectively. The F1 score is the harmonic mean of Precision and Recall, $\text{F1}=\frac{2\,\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$ Note that this F1 measures the cleaner’s quality, not the classifier’s accuracy. , averaged over $10$ runs with $95\%$ confidence intervals). Our zeroth-order method (see Algorithm˜2) adapted to the optimistic formulation attains higher accuracy and F1-score in three different pollute-rates ($p=40\%,60\%,80\%$). Our method’s high performance stems from the minima-selection step: by choosing the iterate with the lowest validation loss along the trajectory, it produces higher-quality cleaner parameters $\theta$, yields better model parameters $x$, and ultimately improves generalization on the held-out test set.
@@
40 RHG $55.64\pm 0.15$ $26.5\pm 0.1$ $54.64\pm 0.19$ $16.5\pm 0.1$
IAPTT-GM $60.4\pm 0.12$ $42.3\pm 0.1$ $59.37\pm 0.15$ $26.5\pm 0.1$
G-PBGD $73.89\pm 0.13$ $50.9\pm 0.1$ $72.54\pm 0.12$ $65.8\pm 0.1$
V-PBGD $65.34\pm 0.13$ $53.2\pm 0.1$ $64.77\pm 0.1$ $55.2\pm 0.1$
PSZO-MinSel $\mathbf{82.85\pm 0.12}$ $\mathbf{70.1\pm 0.1}$ $\mathbf{85.31\pm 0.24}$ $\mathbf{69.1\pm 0.1}$
60 RHG $51.24\pm 0.19$ $23.8\pm 0.1$ $52.94\pm 0.12$ $13.5\pm 0.1$
IAPTT-GM $59.52\pm 0.18$ $23.1\pm 0.1$ $58.72\pm 0.15$ $24.5\pm 0.1$
G-PBGD $64.56\pm 0.12$ $50.2\pm 0.1$ $62.65\pm 0.12$ $51.7\pm 0.1$
V-PBGD $63.14\pm 0.15$ $48.7\pm 0.1$ $62.63\pm 0.13$ $50.4\pm 0.1$
PSZO-MinSel $\mathbf{79.50\pm 0.18}$ $\mathbf{55.2\pm 0.1}$ $\mathbf{80.70\pm 0.18}$ $\mathbf{55.9\pm 0.1}$
80 RHG $48.14\pm 0.12$ $21.5\pm 0.1$ $50.32\pm 0.14$ $11.9\pm 0.1$
IAPTT-GM $56.94\pm 0.14$ $18.1\pm 0.1$ $58.24\pm 0.16$ $17.9\pm 0.1$
G-PBGD $46.84\pm 0.12$ $28.9\pm 0.1$ $47.70\pm 0.14$ $27.3\pm 0.1$
V-PBGD $61.23\pm 0.12$ $33.5\pm 0.1$ $60.42\pm 0.13$ $34.9\pm 0.1$
PSZO-MinSel $\mathbf{69.45\pm 0.15}$ $\mathbf{35.2\pm 0.1}$ $\mathbf{70.27\pm 0.18}$ $\mathbf{44.1\pm 0.1}$
@@ 7 Conclusion and Future Work
This work makes minima selection in bilevel optimization mathematically principled and algorithmically implementable. By working under the condition and exploiting the manifold structure of the lower-level solution set, we introduced a dimension-aware viewpoint: the difficulty of BLO with minima selection is governed by the intrinsic dimension of the minimizer manifold, rather than the ambient dimension of the lower-level variable. The superquantile–Gibbs (SQ-G) relaxation makes this viewpoint concrete, turning minima selection into a controlled sampling-and-selection procedure. More broadly, our analysis shows that minima selection need not be treated as a black box: geometry (via the structure of the minimizer manifold), stochastic approximation (via Gibbs measures), and risk-type functionals (via superquantiles) can be combined to design BLO methods that are both principled and systematically adjustable.
@@ Appendix A Definitions and technical lemmas
This appendix collects the background components used throughout the paper. We recall the Riemannian notions (distance, second fundamental form, injectivity radius, curvature) needed for the geometric arguments in Section˜3. Next, we briefly explain why our standing assumptions imply the technical conditions required to apply [hasenpflug2024wasserstein, Theorem 3.6] in the discussion of Gibbs-measure convergence.

### v4

@@ 6 Experiments
Compute budget and reporting protocol.
Our method is zeroth-order and each outer update invokes a Gibbs–superquantile oracle implemented via PSGD and (parallel-chain) Langevin Monte Carlo, so the wall-clock cost per outer iteration can be higher than that of gradient-based baselines. Accordingly, in the main text we report iteration-based results to illustrate algorithmic behavior and minima-selection effects, and in Appendix F we additionally provide compute-matched comparisons under a fixed wall-clock time budget (including per-iteration timings for all methods).
Experiments are run on a single machine with CPU: Apple M1 Pro (10 cores), RAM: 16 GB with no discrete GPU. Code to reproduce all experiments is released at GitHub. Toy-example methods are implemented in JAX, while data hyper-cleaning methods are implemented in PyTorch.
@@ 6.1 Toy Example
and $u(\theta)=[\cos\theta,\sin\theta,0,\dots,0]^{\top}\in\operatorname{\mathbb{R}}^{d}$. The solution set is the k-dimensional sphere of radius $\sqrt{1+\theta^{2}}$ embedded in $\mathbb{R}^{d}$ along the first $\texttt{k}+1$ coordinates:
@@
    V-PBGD [77]: joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\gamma(g(\theta,x)-g^{*}(\theta))$,
@@
    G-PBGD [77]: joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\tfrac{\gamma}{2}\|\nabla_{x}g(\theta,x)\|^{2}$,
@@
    RHG [29]: truncated reverse‐mode hypergradient with re‐initialization of $\theta$ every outer step,
@@
    IAPTT-GM [62]: unroll $T$ inner steps $x_{k+1}=x_{k}-\beta\nabla_{x}g(\theta,x_{k})$ from $x_{0}=z$, pick the pessimistic truncation $k_{\star}=\arg\max_{k}F(\theta,x_{k})$, and update $\theta,z$ by gradient descent on the truncated objective $F(\theta,x_{k_{\star}})$.
In this part, all methods are implemented in JAX with identical initialization $\theta_{0}=1.0$, $x_{0}\sim\mathcal{N}(0,I_{d})$, and random seed.
@@
Figure 1 plots $F(\theta_{n})-F^{*}$ over iterations for $d=2$. The zeroth-order method rapidly converges toward zero, while the other methods tend to plateau or oscillate around suboptimal objective values, primarily due to the absence of a minima-selection phase.
@@ 6.1.1 Dimension Sweep and Minima‐selection
5 0.273 0.160 0.302 0.151 $\mathbf{0.015\pm 0.012}$
10 0.169 0.153 0.303 0.152 $\mathbf{0.023\pm 0.017}$
20 0.113 0.174 0.302 0.154 $\mathbf{0.043\pm 0.025}$
Table 1: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, where $\hat{\theta}_{N}$ is the selected iterate defined by the best objective value along the trajectory. PSZO-MinSel reports mean $\pm$ 95% CI over 10 runs.
In this experiment, we examine the setting where the intrinsic dimension scales with the ambient dimension, specifically setting $\texttt{k}=d-1$. This simplifies the lower-level objective to $g(\theta,x)=\frac{1}{4}(\|x\|^{2}-\theta^{2})^{2}-\frac{1}{2}\|x\|^{2}$. Table˜1 reports the best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations across varying dimensions $d$, where $\hat{\theta}_{N}$ denotes the selected iterate with the smallest achieved hyper-objective value along the run. Despite this scaling, our algorithm consistently outperforms the three baseline methods, achieving a significantly closer approximation to the global minimizer $\theta^{\ast}=0$. For robustness, we report the mean and confidence intervals over $10$ independent runs for our randomized method, compared to single runs for the deterministic baselines. Baselines are deterministic given initialization; repeated runs match up to numerical tolerance.
@@ 6.1.2 Constant Manifold Dimension
We next investigate the distinction between intrinsic and ambient dimensionality by fixing the manifold dimension at $\texttt{k}=1$ while varying the ambient dimension $d$, using the corresponding instance of $g_{\texttt{k}}(\theta,x)$ as the lower-level objective. To align the implementation with our theoretical bounds, we chose the number of Langevin Monte Carlo (LMC) iterations and the batch size $M$ to closely approximate the target Gibbs distribution. As demonstrated in Table˜2, the best-so-far error of Algorithm˜2 remains nearly constant as $d$ increases, confirming that the complexity is almost independent of the ambient dimension.
@@
5 0.495 0.313 0.302 0.119 $\mathbf{0.015\pm 0.014}$
10 0.499 0.313 0.316 0.114 $\mathbf{0.018\pm 0.012}$
20 0.511 0.315 0.307 0.121 $\mathbf{0.018\pm 0.011}$
30 0.532 0.313 0.311 0.121 $\mathbf{0.02\pm 0.02}$
Table 2: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, with a fixed intrinsic dimension $\texttt{k}=1$.
@@
5 $\mathbf{0.011\pm 0.010}$ $\mathbf{0.021\pm 0.011}$
10 $\mathbf{0.025\pm 0.020}$ $\mathbf{0.023\pm 0.024}$
20 $\mathbf{0.030\pm 0.025}$ $\mathbf{0.026\pm 0.023}$
30 $\mathbf{0.045\pm 0.029}$ $\mathbf{0.027\pm 0.021}$
Table 3: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, evaluated using Algorithm˜2 in the pessimistic setting. The first column corresponds to (33) with lower-level $g_{d-1}$ from (32); the second uses the same problem with lower-level $g_{1}$ from (32).
@@ 6.2 Data Hyper-cleaning
In this section, we evaluate our method on the data hyper-cleaning benchmark [29, 76], a canonical BLO task widely used in the machine learning community. Given a set of corrupted [^20] A sample is clean if its label equals the true class; otherwise it is corrupted. In MNIST, we corrupt by flipping the label of a fraction $p$ of training points uniformly to a different (incorrect) class. $p$ is the pollute-rate. training samples $\mathcal{D}_{\text{tr}}=\{(d_{i}^{\text{tr}},\,\ell_{i}^{\text{tr}})\}_{i=1}^{N}$, a small clean validation set $\mathcal{D}_{\text{val}}=\{(d_{i}^{\text{val}},\,\ell_{i}^{\text{val}})\}_{i=1}^{M}$, and a clean test set, the goal is to learn a data cleaner that down-weights mislabeled training points so as to improve generalization on unseen data. Let $\theta$ denote the data-cleaner parameters (producing per-example weights) and $x$ the model parameters. We use a sigmoid parameterization for the weights, $\omega_{i}(\theta)\;=\;(1+\exp(-\theta_{i}))^{-1},\ (i=1,\dots,N),$ and constrain $\|\theta\|\leq B$ to avoid exploding weights. The task is cast as the optimistic bilevel problem in the following:
@@
We use MNIST with N=5, [^0] N=5{,}000 training, $M=100$ validation, and 10,00010{,}000 test examples, and uniformly corrupt a fraction $p\in\{0.4,0.6,0.8\}$ of the training labels. In practice, the corrupted/imprecise data are widely available, but clean/precise data are scarce. Thus we intentionally adopt a very small validation set. Much of the literature uses roughly equal training/validation splits, in which case the effect of the corruption rate in the uncleaned training data during training is often not apparent. We compare our zeroth-order method (Algorithm˜2) against V-PBGD and G-PBGD [77], RHG [29], and IAPTT-GM [62], using both a linear model and a two-layer MLP which are standard Neural network models in data hyper-cleaning tasks.
@@ Results.
Table˜4 reports solution quality (test accuracy and F1 score [^21] F1-score is defined as follows: We take clean samples as the positive class and corrupted samples as the negative class. Each sample $i$ is assigned a score $s_{i}=\sigma(\theta_{i})$. We predict “clean” for the top (1p)(1-p)-fraction by $s_{i}$. We compute $\text{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\ \text{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$ where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{FN}$ denote true positives, false positives, and false negatives, respectively. The F1 score is the harmonic mean of Precision and Recall, $\text{F1}=\frac{2\,\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$ Note that this F1 measures the cleaner’s quality, not the classifier’s accuracy. , averaged over $10$ runs with $95\%$ confidence intervals). Our zeroth-order method (see Algorithm˜2) adapted to the optimistic formulation attains higher accuracy and F1-score in three different pollute rates ($p=40\%,60\%,80\%$). Our method’s high performance stems from the minima-selection step: by choosing the iterate with the lowest validation loss along the trajectory, it produces higher-quality cleaner parameters $\theta$, yields better model parameters $x$, and ultimately improves generalization on the held-out test set. Appendix F additionally reports compute-matched comparisons under a fixed wall-clock time budget.
@@
40 RHG $55.64\pm 0.45$ $36.5\pm 0.4$ $54.64\pm 0.29$ $36.4\pm 0.2$
IAPTT-GM $60.4\pm 0.22$ $54.3\pm 0.4$ $59.37\pm 0.25$ $52.5\pm 0.4$
G-PBGD $73.89\pm 0.13$ $65.9\pm 0.3$ $72.54\pm 0.32$ $65.8\pm 0.3$
V-PBGD $65.34\pm 0.33$ $59.2\pm 0.2$ $64.77\pm 0.2$ $57.2\pm 0.2$
PSZO-MinSel $\mathbf{82.85\pm 0.32}$ $\mathbf{72.1\pm 0.4}$ $\mathbf{85.31\pm 0.24}$ $\mathbf{73.2\pm 0.5}$
60 RHG $51.24\pm 0.19$ $29.8\pm 0.1$ $52.94\pm 0.42$ $28.5\pm 0.5$
IAPTT-GM $59.52\pm 0.38$ $43.2\pm 0.5$ $58.72\pm 0.35$ $44.2\pm 0.3$
G-PBGD $64.56\pm 0.52$ $50.2\pm 0.5$ $62.65\pm 0.42$ $51.7\pm 0.4$
V-PBGD $63.14\pm 0.43$ $48.7\pm 0.5$ $62.63\pm 0.53$ $50.4\pm 0.4$
PSZO-MinSel $\mathbf{79.50\pm 0.18}$ $\mathbf{55.2\pm 0.1}$ $\mathbf{80.70\pm 0.28}$ $\mathbf{55.9\pm 0.2}$
80 RHG $48.14\pm 0.22$ $21.5\pm 0.2$ $50.32\pm 0.34$ $21.9\pm 0.3$
IAPTT-GM $56.94\pm 0.44$ $27.5\pm 0.3$ $58.24\pm 0.36$ $26.1\pm 0.3$
G-PBGD $46.84\pm 0.32$ $28.9\pm 0.3$ $47.70\pm 0.44$ $27.3\pm 0.3$
V-PBGD $61.23\pm 0.32$ $33.5\pm 0.3$ $60.42\pm 0.42$ $34.9\pm 0.4$
PSZO-MinSel $\mathbf{69.45\pm 0.45}$ $\mathbf{35.2\pm 0.4}$ $\mathbf{70.27\pm 0.48}$ $\mathbf{44.1\pm 0.3}$
@@ 7 Conclusion and Future Work
This paper studies minima-selection in bilevel optimization and proposes a method that is both theoretically grounded and implementable. Under the condition, exploiting the manifold structure of the lower-level solution set, allows to introduce a dimension-aware viewpoint: the difficulty of BLO with minima-selection is governed by the intrinsic dimension of the minimizer manifold, rather than the ambient dimension of the lower-level variable. The proposed superquantile–Gibbs (SQ-G) relaxation turns minima-selection into a sampling-based procedure with an explicit selection rule. Overall, the results suggest that minima-selection can be handled in a structured way by combining geometric information (the minimizer manifold), stochastic approximation (Gibbs measures), and risk-type criteria (superquantiles).
@@ Appendix A Definitions and Technical Lemmas
This appendix collects the background components used throughout the paper. We first fix notation and recall a few basic Riemannian notions for the minimizer manifold (distance, second fundamental form, injectivity radius, and curvature) that are repeatedly used in Section˜3. We then briefly summarize standard differentiability results for bilevel hyper-objectives. Finally, we verify that our assumptions imply the technical conditions required to apply [38, Theorem 3.6] in the discussion of Gibbs-measure convergence.